OpenStudy (anonymous):

hellppp pleasee Determine whether the conjecture is true or false. Give a counterexample for any false conjecture. Given: m2 + 6 = 10 Conjecture: m = 2 is the ONLY solution

OpenStudy (anonymous):

i thought it was true

OpenStudy (anonymous):

do you mean \(m2\) or \(m^2\)?

OpenStudy (anonymous):

\[m ^{2}\]

OpenStudy (anonymous):

then it is false.. there are two solutions possible why.. because the equation is of second degree -> means -> m can be two values \[ m^2+6=10\\ m^2=4 \] now, remember that \[(a)^2=a^2\qquad\text{and}\qquad(-a)^2=a^2\] hence, \[m=\pm\sqrt{4}\] so, m=+2 or m = -2 (two possible solutions which can work)

OpenStudy (anonymous):

thanks